tan^2(x)csc^2(x)-tan^2(x)
1
Expert's answer
2013-03-01T05:48:20-0500
tan^2(x)csc^2(x)-tan^2(x)=tan^2(x) *(csc^2(x)-1)=tan^2(x)*( 1/sin^2(x) -1)=tan^2(x)*( (1-sin^2(x) )/sin^2(x)
)=tan^2(x) * cos^2(x)/sin^2(x)=( sin^2(x)/cos^2(x) ) * cos^2(x)/sin^2(x)=1